LASCAR AND MORLEY RANKS DIFFER IN DIFFERENTIALLY CLOSED FIELDS

EHUD HRUSHOVSKI THOMAS SCANLON

Abstract. We note here, in answer to a question of Poizat, that the Morley and Lascar ranks need not coincide in differentially closed fields. We approach this through the (perhaps) more fundamental issue of the variation of Morley rank in families.